# Nucleon-charmonium interactions from lattice QCD

### v1

@@ 1 Introduction
Experimentally, the low-energy $N$-$c\bar{c}$ interaction, however, is poorly constrained due to the lack of a direct measurement of $N$-$c\bar{c}$ two-body system. For instance, the $N$-$J/\psi$ scattering lengths inferred from analyzing data of the $J/\psi$ photoproduction off the proton with the vector meson dominance and with the low-energy unitarity are $O(1\sim 10)\times 10^{-3}$ fm [26] and $O(1)$ fm [27], respectively. Pioneer lattice studies are limited by quenched approximation [28, 29], the use of heavy pion mass [30], and/or large uncertainties leading to scattering length effectively zero [31, 32].
@@ 5.1 Phase shifts and scattering parameters
where $\rho_{\text{nm}}=0.17~{}\text{fm}^{-3}$ denotes the normal nuclear matter density, and the $N$-$J/\psi$ spin-averaged scattering length is defined as $a^{\text{spin-av}}_{NJ/\psi}=\frac{2a^{\text{(3/2)}}_{0}+a^{\text{1/2)}}_{0}}{3}$.
@@ 6 Summary and conclusions
In summary, we present a realistic study on the low-energy interaction between a nucleon ($N$) and a charmonium ($J/\psi$ and $\eta_{c}$) based on ($2+1$)-flavor lattice QCD calculations with nearly physical light quark masses ($m_{\pi}=146$ MeV) and physical charm quark mass. The $N$-$c\bar{c}$ potential derived from the corresponding hadronic correlation function is found to be attractive at all distances, and its long-range ($1.80.9<r<$ 1.8 fm) part is identified to be consistent with the two-pion exchange interaction. The resulting scattering lengths are $0.30(2)\left({}^{+0}_{-2}\right)$ fm, $0.38(4)\left({}^{+0}_{-3}\right)$ fm, and $0.21(2)\left({}^{+0}_{-1}\right)$ fm for $N$-$J/\psi$ with spin $3/2$, with spin $1/2$, and $N$-$\eta_{c}$, respectively. These results are orders-of-magnitude larger than those inferred from $J/\psi$ photoproduction off the proton assuming the vector meson dominance, and are largely improved compared with previous lattice calculations under the quenched approximation and at heavy pion mass. The results presented here could be checked by future measurements of the femtoscopic correlations of $p$-$J/\psi$.

### v2

@@ 1 Introduction
Experimentally, the low-energy $N$-$c\bar{c}$ interaction, however, is poorly constrained due to the lack of a direct measurement of $N$-$c\bar{c}$ two-body system. For instance, the $N$-$J/\psi$ scattering lengths inferred from analyzing data of the $J/\psi$ photoproduction off the proton with the vector meson dominance and with the low-energy unitarity are $O(1\sim 10)\times 10^{-3}$ fm [26] and $O(1)$ fm [27], respectively. Pioneering lattice studies are limited by quenched approximation [28, 29, 54], the use of heavy pion mass [30], and/or large uncertainties leading to scattering length effectively zero [31, 32].
@@ 5.1 Phase shifts and scattering parameters
where $\rho_{\text{nm}}=0.17~{}\text{fm}^{-3}$ denotes the normal nuclear matter density, and the $N$-$J/\psi$ spin-averaged scattering length is defined as $a^{\text{spin-av}}_{NJ/\psi}=\frac{2a^{\text{(3/2)}}_{0}+a^{\text{(1/2)}}_{0}}{3}$.
@@ 6 Summary and conclusions
In summary, we present a realistic study on the low-energy interaction between a nucleon ($N$) and a charmonium ($J/\psi$ and $\eta_{c}$) based on ($2+1$)-flavor lattice QCD calculations with nearly physical light quark masses ($m_{\pi}=146$ MeV) and physical charm quark mass. The $N$-$c\bar{c}$ potential derived from the corresponding hadronic correlation function is found to be attractive at all distances, and its long-range ($1.80.9<r<$ 1.8 fm) part is identified to be consistent with the two-pion exchange interaction. The resulting scattering lengths are $0.30(2)\left({}^{+0}_{-2}\right)$ fm, $0.38(4)\left({}^{+0}_{-3}\right)$ fm, and $0.21(2)\left({}^{+0}_{-1}\right)$ fm for $N$-$J/\psi$ with spin $3/2$, with spin $1/2$, and $N$-$\eta_{c}$ with spin $1/2$, respectively. These results are orders-of-magnitude larger than those inferred from $J/\psi$ photoproduction off the proton assuming the vector meson dominance, and are largely improved compared with previous lattice calculations under the quenched approximation and at heavy pion mass. The results presented here could be checked by future measurements of the femtoscopic correlations of $p$-$J/\psi$, which have been shown to be sensitive to the underlying interaction [63, 64].